OpenStudy (anonymous):

How many 100W light bulbs could be powered for one year by the direct conversion of 1 g of matter into energy?

OpenStudy (anonymous):

E=mc^2=.001*(3*10^8)^2=.009*10^16=9*10^13 joule 100 t=9*10^13 t=9*10^11 now convert this seconds into year....i think u can do it.....

OpenStudy (anonymous):

E=mc^2=.001*(3*10^8)^2=.009*10^16=9*10^13 joule now let the number of bulbs be n,so n*100* t=9*10^13 t=365*3600 substitue to find n, the number of bulbs u can glow.

OpenStudy (anonymous):

now let the number of bulbs be n,so n*100* t=9*10^13 t=365*3600 substitue to find n, the number of bulbs u can glow. ^ What formula are you applying here? I am confused as to what you did here.